OpenStudy (anonymous):

How many rectangles (including squares) are on an 8x8 chessboard? Bonus imaginary points for the general form for any size chessboard.

OpenStudy (anonymous):

(Thinking Mathematically, page 39)

OpenStudy (dumbcow):

is it 263? i'll skip on the imaginary points

OpenStudy (anonymous):

It is not.

OpenStudy (dumbcow):

261?

OpenStudy (anonymous):

Nopes.

OpenStudy (dumbcow):

oh well you stumped me on this one :)

OpenStudy (anonymous):

Talk me through your reasoning :)

OpenStudy (akshay_budhkar):

1296

OpenStudy (dumbcow):

i went through all the 1Xn rectangles, then all the 2Xn, up to 8Xn where n is from 1 to 8

OpenStudy (akshay_budhkar):

???????/

OpenStudy (anonymous):

How did you figure out how many 1xn rectangles there were? Ashkay, you are quite correct!

OpenStudy (dumbcow):

64 1X1 32 1X2 20 1X3 16 1X4 ....

OpenStudy (dumbcow):

oh wait i didn't allow for overlapping ..haha

OpenStudy (anonymous):

Aha, there are 56 1x2 rectangles due to overlapping :P

OpenStudy (anonymous):

There are still bonus points going for the general form for any size chessboard. I'll post the answer in about 30 minutes if people want to try to get it!

OpenStudy (akshay_budhkar):

actually it was a guess... i was getting something around 1200 but i rounded it to 1296 as it is generally the answer:)

OpenStudy (anonymous):

Hah, well played sir.

OpenStudy (akshay_budhkar):

i got 204 squares

OpenStudy (akshay_budhkar):

still i want the solution.... i am trying again though for perfect answer

OpenStudy (akshay_budhkar):

got it. thanks to my dad

OpenStudy (akshay_budhkar):

it is [(n)(n+1)/2]^2....

OpenStudy (anonymous):

What's n?

OpenStudy (akshay_budhkar):

n*n chess board

OpenStudy (akshay_budhkar):

right?

OpenStudy (anonymous):

It's a different formula to the one I have, but it seems to be correct!

OpenStudy (akshay_budhkar):

whats yours?

OpenStudy (anonymous):

Aha, I see how they link together now. Mine is \[\sum_{n=1}^{8}\sum_{m=1}^{8}(9-m)(9-n)\]

OpenStudy (akshay_budhkar):

its the same!!!!

OpenStudy (anonymous):

Or the more general form for any size chessboard (dimensions x and y) is \[\sum_{n=1}^{y}\sum_{m=1}^{x}((x+1)-m)((y+1)-n)\]

OpenStudy (akshay_budhkar):

how did you get it?

OpenStudy (anonymous):

m and n are the dimensions of the rectangles.

OpenStudy (anonymous):

Take a rectangle of width 1, and height 1, and place it on the chessboard. You'll be able to fit 8 of the rectangles across the board, and 8 rectangles down the board. The total of those rectangles that will fit on the board is the product of those: 64. Move onto another size rectangle mxn. You can fit \((9-m)\) rectangles across the board, and \((9-n)\) rectangles down the board. The total number of these rectangles that can fit on the board is the product: \((9-m)(9-n)\). To take into account all values of m, and n, you have to do a summation where m, and n can be 1-8.

OpenStudy (anonymous):

Summations starting at 1 go as \(\frac{(n)(n+1)}{2}\), which is what your dad presumably figured out.

OpenStudy (anonymous):

The square comes into it because the board is square I thin. If the board has dimensions \(x\times y\), then it would be more like \[ \frac{(x)(x+1)}{2} \times \frac{(y)(y+1)}{2} \]